# Proxy-GS: Efficient 3D Gaussian Splatting via Proxy Mesh

### v2

@@ 5 Experiment
We begin by comparing our approach with other methods on the large-scale urban dataset (Li et al., 2023) to assess rendering quality. We follow the partition script of the MatrixCity, and divided the 8477 street images in its Small City into 5 blocks. Details can be seen in the Appendix A.1. The evaluation is further extended to large-scale indoor scenes from Zip-NeRF (Barron et al., 2023). In addition, we also test on real-world street scenes from the Small City dataset (Kerbl et al., 2024), as well as real-world aerial-view scenes from CUHK-LOWER (Xiong et al., 2024), which contain relatively fewer occlusions.
@@
We adopt three widely used image quality metrics to evaluate novel view synthesis: peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018). In addition, we report frames per second (FPS) to measure the rendering efficiency of different methods.
Figure 4: Qualitative comparison. Visualization on different datasets (Li et al., 2023; Xiong et al., 2024; Barron et al., 2023; Kerbl et al., 2024).
@@
Our method is implemented on top of the state-of-the-art MLP-based Octree-GS (Ren et al., 2024), following its default initialization and LOD strategy. For comparison, we also re-implement 3DGS (Kerbl et al., 2023a), Scaffold-GS (Lu et al., 2024), and Hierarchical-GS (Kerbl et al., 2024), and train all methods for 40k iterations. Specifically, for the evaluation of Hierarchical-GS, we set the $\tau_{1},\tau_{2},\tau_{3}=3,6,15$. For approaches that do not employ MLPs, such as 3DGS and Hierarchical-GS, their default configurations typically yield higher rendering FPS but exhibit a noticeable quality gap compared to Octree-GS. Since an increased number of Gaussian primitives generally leads to better rendering quality (Zhao et al., 2024), we reduce the densification threshold to $10^{-4}$ across all scenes to ensure a fair comparison, resulting in rendering quality closer to that of Octree-GS. Unlike Octree-GS, Scaffold-GS initializes with fewer anchors due to the absence of multi-round sampling. To improve its rendering fidelity, we adopt a smaller voxel size of $10^{-4}$ together with a lower densification threshold of $10^{-4}$. All training experiments are performed on a single NVIDIA A100-40GB GPU. For inference, we employ a consumer-grade RTX 4090 GPU to reflect real-world deployment scenarios better.
@@ 5.1 Main Results
Furthermore, to demonstrate the generality of our method, in Tab. 2 We also evaluate our method on aerial-view scenes, indoor environments, and real-world town streets, where it achieves comparable or superior performance against current state-of-the-art methods. Although aerial scenes typically involve limited occlusions and the current indoor dataset often contains relatively few rooms with sparse occlusion patterns, our method still yields noticeable improvements. Moreover, for small-city street scenes, which bear resemblance to the MatrixCity dataset, our approach delivers substantial improvements over the MLP-based method Octree-GS, achieving higher rendering quality while boosting FPS by nearly $3\times$. These results collectively demonstrate the broad applicability of our method across diverse scenarios, while also highlighting that the extent of performance gains may vary depending on the characteristics of the scene.
@@ 5.2 Ablations
Table 3: Ablations of different safety margin of depth culling $\gamma$ trained on Small City Kerbl et al. (2024).
@@
0.1 22.94 0.734 0.349 142
0.3 23.09 0.736 0.344 139
0.6 23.02 0.735 0.348 135
1.0 23.05 0.736 0.345 128
ID 3 further enforces consistency by employing proxy-guided rendering also during training. In this case, rendering quality surpasses the baseline, while FPS slightly decreases compared to ID 2, mainly because more anchors grow before being culled by occlusion.
ID 4 incorporates the proposed proxy-guided densification strategy in addition to proxy-guided training and rendering. This setting achieves the best balance, delivering further improvements in rendering quality while maintaining a comparable FPS to ID 3.
ID
Occlusion
Training
Proxy-guided
Densification
Proxy-guided
Inference
[…]
Average
anchor
@@
Figure 5: Visualization on different safety margins.
Safety margin of the occlusion culling.
In the Appendix A.2, we report all the results with the hardware 3DGS as the default renderer
@@
In Fig. 3, we quantify the proportion of inference time spent on each component. The lightweight proxy-based depth rendering takes nearly negligible time (around $1\text{\,}\mathrm{m}\mathrm{s}$). Our anchor filtering is also faster due to the reduced number of anchors. The rendering stage is where most of the savings come from: with fewer anchors, both the decoding overhead and Gaussian rasterization are significantly reduced. For more details, we also record the average decode anchors in the Appendix A.3.
Table 5: Integration with different 3DGS rendering accelerations. We evaluate our method combined with existing approaches (Feng et al., 2024; fas, 2024) on Block 1.
@@
Since our method primarily optimizes anchors and thus indirectly reduces the number of rendered Gaussians, it can be naturally combined with existing acceleration techniques for the original 3DGS to achieve even higher speed. In Table 5, we evaluate on Block 1. Here, Original 3DGS denotes the default renderer used in Proxy-GS. Replacing it with FlashGS brings a minor improvement, while using a hardware rasterizer for 3DGS slightly compromises rendering quality but further boosts the frame rate by nearly 40 FPS. .
@@ 6 Conclusion
Reproducibility
To ensure reproducibility, we will release the complete training and inference code. All results reported in this paper can be reproduced using the released repository, along with the same experimental settings described in the main text and Appendix. The final model of all the datasets will also be made publicly available.
Ethics Statement
Our research is devoted to enhancing the efficiency and rendering quality of 3D reconstruction techniques. No experiments involve human participants, personally identifiable information, or sensitive content. The datasets employed are openly released for academic purposes and have been broadly utilized in prior literature. We adhered to their licensing terms and conducted all experiments in a manner consistent with data privacy and integrity. We consider this work to present no evident ethical concerns or potential societal risks.
Appendix A Appendix
A.1 Division detail on MatrixCity
We divide all the Horizon street scenes in MatrixCity’s small city into five blocks (eg. Block 1, Block 2), The partition margin details is in Tab.  6
A.2 Combine with Hardware 3DGS
We combine our method with Hardware 3DGS (fas, 2024) in Tab. 7 and Tab. 8. As observed, the FPS improves across all datasets, but due to the precision settings used, there is a noticeable decline in rendering quality.
Table 6: Partition information in MatirxiCity.
@@ A.2 Combine with Hardware 3DGS
Table 7: Combine with Hardware 3DGS  (fas, 2024), quantitative results on MatrixCity (Li et al., 2023)
@@
Table 8: Combine with Hardware 3DGS  (fas, 2024), quantitative results on real world Outdoor and Indoor datasets (Xiong et al., 2024; Kerbl et al., 2024; Barron et al., 2023).
@@
A.3 Average decoded anchor number on all the datasets
@@
A.4 Mesh extraction on different datasets
A.4.1 Indoor and Outdoor Scenes with Dense Point Clouds
We describe the mesh extraction process when dense point clouds are available for both indoor and outdoor environments. This category includes real-world datasets that provide LiDAR point clouds (e.g., Xiong et al. (2024)), where mesh generation can be directly performed using surface reconstruction methods, such as  (Huang et al., 2023). In addition, for synthetic datasets such as MatrixCity, ground-truth depth maps are available, which can be fused via TSDF to obtain high-quality meshes.
A.4.2 Indoor Scenes with Sparse COLMAP Point Clouds
We describe the workflow of mesh extraction in indoor scenes where only sparse COLMAP reconstructions are available. Directly relying on COLMAP to generate dense point clouds in indoor environments is often unreliable, as such scenes frequently contain large textureless regions. To address this challenge, we follow the recent advances in 3DGS-based indoor reconstruction methods (Ren et al., 2025; Zhang et al., 2024), where sparse texture cues are complemented by state-of-the-art monocular depth estimation (e.g., MoGe2 (Wang et al., 2025)) and further refined with multi-view constraints from PGSR Chen et al. (2024). This hybrid strategy enables the recovery of reasonable indoor meshes despite the limitations of sparse COLMAP input.
A.4.3 Outdoor Scenes with Sparse COLMAP Point Clouds
For outdoor environments where the reconstruction relies solely on sparse COLMAP point clouds, the abundance of feature points generally mitigates the issue of sparse textures. However, due to the large spatial extent, many 3DGS-based indoor reconstruction methods encounter out-of-memory (OOM) problems when applied to outdoor scenes. To address this, we employ CityGS-X (Gao et al., 2025), a state-of-the-art large-scale geometric reconstruction framework, which leverages multi-GPU parallelism to achieve scalable mesh generation with competitive performance.
A.4.4 Mesh visualization
@@
Figure 6: Mesh visualization. Scenes include different datasets (Li et al., 2023; Xiong et al., 2024; Barron et al., 2023; Kerbl et al., 2024).
A.5 Fast Depth Acquisition
A.5.1 Overview.
@@ A.5 Fast Depth Acquisition
A.5.2 Preprocessing: from reconstructed mesh to clusters.
@@
    $Q\;=\;\sum_{f}\mathbf{p}_{f}\mathbf{p}_{f}^{\top},\qquad$
@@
    $Q^{\prime}\;=\;Q_{i}+Q_{j},\qquad E(\tilde{\mathbf{x}})\;=\;\tilde{\mathbf{x}}^{\top}Q^{\prime}\tilde{\mathbf{x}}$.
@@
    $\mathbf{x}^{*}\;=\;\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{3}}\;\mathbf{x}^{\top}A\,\mathbf{x}+2\,\mathbf{b}^{\top}\mathbf{x}+c\;=\;-\,A^{-1}\mathbf{b}\quad(\text{if }A\text{ is invertible}),$
@@
    $\mathbf{p}_{1}=(\hat{\mathbf{n}},\ -\hat{\mathbf{n}}^{\top}\mathbf{x}_{0})^{\top},\quad\mathbf{p}_{2}=(\widehat{\mathbf{t}\times\hat{\mathbf{n}}},\ -\,\widehat{\mathbf{t}\times\hat{\mathbf{n}}}^{\!\top}\mathbf{x}_{0})^{\top},$
@@
    $\mathbf{y}_{k,j}=PV\!\begin{bmatrix}\mathbf{x}_{k,j}\\ 1\end{bmatrix},\qquad\mathbf{u}_{k,j}^{\text{ndc}}=\Big(\tfrac{y^{x}_{k,j}}{y^{w}_{k,j}},\,\tfrac{y^{y}_{k,j}}{y^{w}_{k,j}}\Big).$
@@
    $R_{k}^{(0)}=\big[\ \lfloor\min_{j}\mathbf{s}_{k,j}\rfloor-\Delta,\ \ \lceil\max_{j}\mathbf{s}_{k,j}\rceil+\Delta\ \big]\ \cap\ [0,W\!-\!1]\times[0,H\!-\!1].$
@@
In order to obtain the depth every frame efficiently, a naive path would be to read back the GPU depth buffer to host memory and then upload it to CUDA, introducing synchronization and PCIe traffic. Instead, we adopt a fully GPU-resident path: we render with Vulkan and export the depth image’s memory as an external file descriptor (FD). On the CUDA side, we import that FD as external memory and map it to a device pointer; the pointer is then wrapped as a PyTorch CUDA tensor without a copy. This eliminates CPU involvement, avoids extra copies, and preserves real-time throughput.
LLM Usage
We acknowledge the assistance of OpenAI’s ChatGPT in the preparation of this manuscript. ChatGPT was used for language refinement, LaTeX formatting, and figure illustration to improve the clarity of the presentation. All conceptual ideas, methodological designs, and final decisions were solely made by the authors.

### v3

@@ 5 Experiment
We begin by comparing our approach with other methods on the large-scale urban dataset [21] to assess rendering quality. We follow the partition script of the MatrixCity, and divided the 8477 street images in its Small City into 5 blocks. Details can be seen in the Appendix. The evaluation is further extended to indoor scenes from Zip-NeRF [2]. In addition, we also test on real-world scenes that have different levels of occlusion and scale, including a street scene: Small City dataset [16], and aerial-view scenes from CUHK-LOWER [34]. We select both indoor and aerial top-down scenes to examine whether our method incurs additional overhead in relatively small or minimally occluded environments.
@@
We adopt three widely used image quality metrics to evaluate novel view synthesis: peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and learned perceptual image patch similarity (LPIPS) [40]. In addition, we report frames per second (FPS) to measure the rendering speed of different methods.
Figure 4: Qualitative comparison. Visualization on different datasets, where the regions with noticeable differences are highlighted and zoomed in with red boxes [21, 34, 2, 16].
@@
Our method is implemented on top of the state-of-the-art MLP-based Octree-GS [30], following its default initialization and LOD strategy. For comparison, we also re-implement 3DGS [14], Scaffold-GS [27], and Hierarchical-GS [16], and train all methods for 40k iterations. Specifically, for the evaluation of Hierarchical-GS, we set the $\tau_{1},\tau_{2},\tau_{3}=3,6,15$. For approaches that do not employ MLPs, such as 3DGS and Hierarchical-GS, their default configurations typically yield higher rendering FPS but exhibit a noticeable quality gap compared to Octree-GS. Since an increased number of Gaussian primitives generally leads to better rendering quality [41], we reduce the densification threshold to $10^{-4}$ across all scenes to ensure a fair comparison, resulting in rendering quality closer to that of Octree-GS. Unlike Octree-GS, Scaffold-GS initializes with fewer anchors due to the absence of multi-round sampling. To improve its rendering fidelity, we adopt a smaller voxel size of $10^{-4}$ together with a lower densification threshold of $10^{-4}$. All training experiments are performed on a single NVIDIA A100-40GB GPU. For inference, we employ a consumer-grade RTX 4090 GPU to reflect real-world deployment scenarios better.
@@ 5.1 Main Results
Furthermore, to demonstrate the generality of our method, in Tab. 2 We also evaluate our method on aerial-view scenes, indoor environments, and real-world town streets, where it achieves comparable or superior performance against current state-of-the-art methods. Although aerial scenes typically involve limited occlusions and the current indoor dataset often contains relatively few rooms with sparse occlusion patterns, our method still yields noticeable improvements. Moreover, for small-city street scenes, which bear resemblance to the MatrixCity dataset, our approach delivers substantial improvements over the MLP-based method Octree-GS, achieving higher rendering quality while boosting FPS by $2.73\times$. These results collectively demonstrate the broad applicability of our method across diverse scenarios, while also highlighting that the extent of performance gains may vary depending on the characteristics of the scene.
Figure 5: Quantitative mesh visualization on different Resolutions and Vertex noise on Block 5.
Figure 6: Quantitative visualization of different Vertex noise on Block 5.
@@ 5.2 Ablations
ID 3 further enforces consistency by employing proxy-guided rendering also during training. In this case, rendering quality surpasses the baseline, while FPS slightly decreases compared to ID 2, mainly because more anchors grow before being culled by occlusion. ID 4 incorporates the proposed proxy-guided densification strategy and proxy-guided training and rendering. This setting achieves the best balance, delivering further improvements in rendering quality while maintaining a comparable FPS to ID 3.
@@
In Fig. 3, we quantify the proportion of inference time spent on each component. The lightweight proxy-based depth rendering takes nearly negligible time (around $1\text{\,}\mathrm{m}\mathrm{s}$). Our anchor filtering is also faster due to the reduced number of anchors. The rendering stage is where most of the savings come from: with fewer anchors, both the decoding overhead and Gaussian rasterization are significantly reduced. For more details, we also record the average decode anchors in the Appendix 7.4.
Table 4: Integration with different 3DGS rendering accelerations. We evaluate our method combined with existing approaches [7, 6] on Block 1.
@@
Since our method primarily optimizes anchors and thus indirectly reduces the number of rendered Gaussians, it can be naturally combined with existing acceleration techniques for the original 3DGS to achieve even higher speed. In Table 4, we evaluate on Block 1. Here, Original 3DGS denotes the default renderer used in Proxy-GS. Replacing it with FlashGS brings a minor improvement, while using a hardware rasterizer for 3DGS slightly compromises rendering quality but further boosts the frame rate by nearly 40 FPS. Note that we reported results with original 3DGS renderer in Table 123. In the Appendix  7.2, we report all the results with the hardware 3DGS as the default renderer
Figure 7: Ablation on proxy fidelity: (a) PSNR vs. Proxy Resolution; (b) PSNR vs. Vertex Noise.
Dependency on different Proxy Quality.
To quantify our method’s dependence on the proxy’s accuracy, we conduct two ablations: (i) mesh resolution, we evaluate proxies at multiple resolutions (from fine (108MB) to coarse (824KB)) ); and (ii) vertex noise, we perturb mesh vertices with random noise of varying magnitudes. As shown in Fig. 7, varying the proxy resolution has only a marginal effect on rendering quality. This is largely because modern urban scenes (buildings, facades, and roads) are dominated by broad, near-planar surfaces, so even coarse proxies preserve the visibility structure needed by our filter. The mesh visualizations in Fig.  5 illustrate this effect clearly. When the mesh resolution is reduced, the overall occlusion structure remains correct, only fine details are lost, so the impact on rendering quality is minimal. In contrast, adding vertex noise disrupts the global geometry and breaks the occlusion boundaries, which inherently leads to a much larger degradation in rendering quality. As illustrated in Fig. 6, increasing noise introduces spurious protrusions on the ground, which smear the image and progressively degrade rendering quality. However, As shown in Fig. 2, because there is an inherent offset between the anchors and the decoded Gaussians, our method retains robustness under small perturbations: with noise levels within 5%, the overall impact remains limited.
@@ 6 Conclusion
Supplementary Material
7 Appendix
7.1 Division detail on MatrixCity
We divide all the Horizon street scenes in MatrixCity’s small city into five blocks (eg. Block 1, Block 2), The partition margin details is in Tab.  5
7.2 Combine with Hardware 3DGS
We combine our method with Hardware 3DGS [6] in Tab. 7 and Tab. 8. As observed, the FPS improves across all datasets, but due to the precision settings used, there is a noticeable decline in rendering quality.
Table 5: Partition information in MatrixCity.
@@ 7.2 Combine with Hardware 3DGS
7.3 Comparison with different depth-acquisition approaches
In Tab. 6, To demonstrate the effectiveness of our depth acquisition strategy, we also evaluate two alternative approaches: directly rendering depth using nvdiffrast [18], and extracting depth from a pre-trained 3DGS model. These two depth-generation baselines allow us to compare our method against commonly used depth sources.
Table 6: The FPS of different depth acquisition methods on Block 5.
Method Nvdiffrast 3DGS Ours
FPS 32 54 151
Table 7: Combine with Hardware 3DGS  [6], quantitative results on MatrixCity [21]
@@ 7.3 Comparison with different depth-acquisition approaches
Table 8: Combine with Hardware 3DGS  [6], quantitative results on real world Outdoor and Indoor datasets [34, 16, 2].
@@
7.4 Average decoded anchor number on all the datasets
@@
Figure 8: Visualization on different safety margins.
Table 10: Ablations of different safety margin of depth culling $\gamma$ trained on Small City [16].
[…]
0.1 22.94 0.734 0.349 142
0.3 23.09 0.736 0.344 139
0.6 23.02 0.735 0.348 135
1.0 23.05 0.736 0.345 128
Safety margin of the occlusion culling.
7.5 Mesh extraction on different datasets
7.5.1 Indoor and Outdoor Scenes with Dense Point Clouds
We describe the mesh extraction process when dense point clouds are available for both indoor and outdoor environments. This category includes real-world datasets that provide LiDAR point clouds (e.g., [34]), where mesh generation can be directly performed using surface reconstruction methods, such as  [11]. In addition, for synthetic datasets such as MatrixCity, ground-truth depth maps are available, which can be fused via TSDF to obtain high-quality meshes.
7.5.2 Indoor Scenes with Sparse COLMAP Point Clouds
We describe the workflow of mesh extraction in indoor scenes where only sparse COLMAP reconstructions are available. Directly relying on COLMAP to generate dense point clouds in indoor environments is often unreliable, as such scenes frequently contain large textureless regions. To address this challenge, To address this challenge, we leverage a foundation-model approach similar to VGGT [32]. Our experiments show that VGGT-style models exhibit strong robustness in indoor scenes. However, since our pipeline requires alignment with COLMAP poses, we adopt MapAnything [13] instead, using RGB images together with COLMAP inputs to obtain a dense point cloud. This hybrid strategy enables the recovery of reasonable indoor meshes despite the limitations of sparse COLMAP input.
7.5.3 Outdoor Scenes with Sparse COLMAP Point Clouds
For outdoor environments where the reconstruction relies solely on sparse COLMAP point clouds, the abundance of feature points generally mitigates the issue of sparse textures. However, due to the large spatial extent, many 3DGS-based indoor reconstruction methods encounter out-of-memory (OOM) problems when applied to outdoor scenes. To address this, we employ CityGS-X [9], a state-of-the-art large-scale geometric reconstruction framework, which leverages multi-GPU parallelism to achieve scalable mesh generation with competitive performance.
7.5.4 Mesh visualization
@@
Figure 9: Mesh visualization. Scenes include different datasets [21, 34, 2, 16].
7.6 Fast Depth Acquisition
7.6.1 Overview.
@@ 7.6 Fast Depth Acquisition
7.6.2 Preprocessing: from reconstructed mesh to clusters.
@@
    $Q\;=\;\sum_{f}\mathbf{p}_{f}\mathbf{p}_{f}^{\top},\hskip 28.80008pt$
@@
    $Q^{\prime}\;=\;Q_{i}+Q_{j},\hskip 28.80008ptE(\tilde{\mathbf{x}})\;=\;\tilde{\mathbf{x}}^{\top}Q^{\prime}\tilde{\mathbf{x}}$.
@@
    $\mathbf{x}^{*}$ $=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{3}}\;\mathbf{x}^{\top}A\,\mathbf{x}+2\,\mathbf{b}^{\top}\mathbf{x}+c$ (7)
    $=-\,A^{-1}\mathbf{b}\quad\text{(if $A$ is invertible).}$
@@
    $\mathbf{p}_{1}=(\hat{\mathbf{n}},\ -\hat{\mathbf{n}}^{\top}\mathbf{x}_{0})^{\top},\qquad\mathbf{p}_{2}=(\widehat{\mathbf{t}\times\hat{\mathbf{n}}},\ -\,\widehat{\mathbf{t}\times\hat{\mathbf{n}}}^{\!\top}\mathbf{x}_{0})^{\top},$
@@
    $\mathbf{y}_{k,j}=PV\!\begin{bmatrix}\mathbf{x}_{k,j}\\ 1\end{bmatrix},\hskip 28.80008pt\mathbf{u}_{k,j}^{\text{ndc}}=\Big(\tfrac{y^{x}_{k,j}}{y^{w}_{k,j}},\,\tfrac{y^{y}_{k,j}}{y^{w}_{k,j}}\Big).$
@@
    $R_{k}^{(0)}$ $=\Bigl[\,\lfloor\min_{j}\mathbf{s}_{k,j}\rfloor-\Delta,\ \ \lceil\max_{j}\mathbf{s}_{k,j}\rceil+\Delta\,\Bigr]$ (8)
    $\qquad\cap\ [0,W\!-\!1]\times[0,H\!-\!1].$
@@
In order to obtain the depth every frame efficiently, a naive path would be to read back the GPU depth buffer to host memory and then upload it to CUDA, introducing synchronization and PCIe traffic. Instead, we adopt a fully GPU-resident path: we render with Vulkan and export the depth image’s memory as an external file descriptor (FD). On the CUDA side, we import that FD as external memory and map it to a device pointer; the pointer is then wrapped as a PyTorch CUDA tensor without a copy.